OpenStudy (anonymous):

Which of the following is a proportion? twenty-eight over four equals fourteen over two seven over fifty-six equals four over twenty-eight four over seven equals three over six twenty-eight over seven equals one over four

Directrix (directrix):

twenty-eight over four equals fourteen over two --> IS a proportion 28/4 = 14/2 7 = 7 seven over fifty-six equals four over twenty-eight --> is NOT a proportion 7/56 = 4/28 1/8 = 1/4 four over seven equals three over six ---> is NOT a proportion 4/7 = 3/6 4/7 = 1/2 twenty-eight over seven equals one over four ---> Is NOT a proportion 28/7= 1/4 4/1 = 1/4